OpenStudy (anonymous):

Which is an equation of a circle with center (5, 0) that passes through the point (1, 1)? (x - 1)^2 + (y - 1)^2 = 17 (x - 5)^2 + y^2 = 17 (x + 5)^2 + y^2 = 17 (x + 5)^2 + y^2 = 24

OpenStudy (agent0smith):

Center (5,0) means the equation has to look like this: (x-5)^2 + y^2 =

OpenStudy (anonymous):

17

OpenStudy (agent0smith):

So... you can already eliminate three answers and don't need to use the point (1,1) at all!

OpenStudy (agent0smith):

If you did though, you'd plug in x=1 and y=1: (1-5)^2 + 1^2 = 17 (-4)^2 + 1 = 17 16 + 1 = 17 which is true :)

OpenStudy (anonymous):

yay ok thank you :)

OpenStudy (agent0smith):

Welcome emuhleeeee!